12-n, 12, 12+n
What is the average (arithmetic mean) of the 3 quantities in the list above?

Found 2 solutions by user_dude2008, JBarnum:
Answer by user_dude2008(1861)

(12-n+12+12+n)/3=36/3=12

Answer by JBarnum(2044)